Quadratic Equations
7. A basketball is tossed from the top of a 3-m wall. The path of the basketball is defined by the relation y=-x^2+2x+3, where x represents the horizontal distance travelled, in metres, and y represents the height, in metres, above the ground. How far has the basketball travelled horizontally when it lands on the ground?
Thank you very much!!!

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

A basketball is tossed from the top of a 3-m wall. The path of the basketball is defined by the relation y=-x^2+2x+3, where x represents the horizontal distance traveled, in metres, and y represents the height, in metres, above the ground. How far has the basketball traveled horizontally when it lands on the ground?
-----------------------------------
The height of the ball is zero when it hits the ground.
---------------------
Solve -x^2 + 2x + 3 = 0
---
(-x + 3)(x+1) = 0
---
Positive solution:
x = 3 meters (horizontal travel distance)
===============================================
Cheers,
Stan H.


Answer by solver91311(24713)


The ground is zero meters above the ground, so you want to know the value of when . Hence:



Just solve the quadratic for .

John